OpenStudy (anonymous):

Simplify using your knowledge of properties. 2(b - 6) + -2b + 12.

OpenStudy (anonymous):

2(b - 6) + -2b + 12 = 2(b) - 2(6) -2b +12 = 2b - 2b -12 + 12 = 0